\begin{thebibliography}{10}

\bibitem{poyatos1991paralinguistic}
Fernando Poyatos.
\newblock Paralinguistic qualifiers: Our many voices.
\newblock {\em Language \& Communication}, 11(3):181--195, 1991.

\bibitem{eyben2010opensmile}
Florian Eyben, Martin W{\"o}llmer, and Bj{\"o}rn Schuller.
\newblock Opensmile: the munich versatile and fast open-source audio feature
  extractor.
\newblock In {\em Proceedings of the 18th ACM international conference on
  Multimedia}, pages 1459--1462. ACM, 2010.

\bibitem{keating2015acoustic}
Patricia Keating, Marc Garellek, and Jody Kreiman.
\newblock Acoustic properties of different kinds of creaky voice.
\newblock In {\em Proceedings of the 18th International Congress of Phonetic
  Sciences}, pages 0821--1, 2015.

\bibitem{doi:10.1044/jshr.3302.298}
L.~Eskenazi, D.~G. Childers, and D.~M. Hicks.
\newblock Acoustic correlates of vocal quality.
\newblock {\em Journal of Speech, Language, and Hearing Research},
  33(2):298--306, 1990.

\bibitem{de1995some}
Guus de~Krom.
\newblock Some spectral correlates of pathological breathy and rough voice
  quality for different types of vowel fragments.
\newblock {\em Journal of Speech, Language, and Hearing Research},
  38(4):794--811, 1995.

\bibitem{childers1991vocal}
Donald~G Childers and CK~Lee.
\newblock Vocal quality factors: Analysis, synthesis, and perception.
\newblock {\em the Journal of the Acoustical Society of America},
  90(5):2394--2410, 1991.

\bibitem{wolfe1987prediction}
Virginia~I Wolfe and Thomas~M Steinfatt.
\newblock Prediction of vocal severity within and across voice types.
\newblock {\em Journal of Speech, Language, and Hearing Research},
  30(2):230--240, 1987.

\bibitem{stranik2014acoustic}
Adam Str{\'a}n{\'\i}k, Roman {\v{C}}mejla, and Jan Vok{\v{r}}{\'a}l.
\newblock Acoustic parameters for classification of breathiness in continuous
  speech according to the grbas scale.
\newblock {\em Journal of Voice}, 28(5):653--e9, 2014.

\bibitem{labuschagne2016perception}
Ilse~Bernadette Labuschagne and Valter Ciocca.
\newblock The perception of breathiness: Acoustic correlates and the influence
  of methodological factors.
\newblock {\em Acoustical Science and Technology}, 37(5):191--201, 2016.

\bibitem{hammarberg1980perceptual}
Britta Hammarberg, Bernard Fritzell, J~Gaufin, Johan Sundberg, and Lage Wedin.
\newblock Perceptual and acoustic correlates of abnormal voice qualities.
\newblock {\em Acta oto-laryngologica}, 90(1-6):441--451, 1980.

\bibitem{hillenbrand1994acoustic}
James Hillenbrand, Ronald~A Cleveland, and Robert~L Erickson.
\newblock Acoustic correlates of breathy vocal quality.
\newblock {\em Journal of Speech, Language, and Hearing Research},
  37(4):769--778, 1994.

\bibitem{klatt1990analysis}
Dennis~H Klatt and Laura~C Klatt.
\newblock Analysis, synthesis, and perception of voice quality variations among
  female and male talkers.
\newblock {\em the Journal of the Acoustical Society of America},
  87(2):820--857, 1990.

\bibitem{hillenbrand1996acoustic}
James Hillenbrand and Robert~A Houde.
\newblock Acoustic correlates of breathy vocal quality: dysphonic voices and
  continuous speech.
\newblock {\em Journal of Speech, Language, and Hearing Research},
  39(2):311--321, 1996.

\bibitem{garellek2012timing}
Marc Garellek.
\newblock The timing and sequencing of coarticulated non-modal phonation in
  english and white hmong.
\newblock {\em Journal of Phonetics}, 40(1):152--161, 2012.

\bibitem{gerfen2005production}
Chip Gerfen and Kirk Baker.
\newblock The production and perception of laryngealized vowels in coatzospan
  mixtec.
\newblock {\em Journal of Phonetics}, 33(3):311--334, 2005.

\bibitem{garellek2015perception}
Marc Garellek.
\newblock Perception of glottalization and phrase-final creak.
\newblock {\em The Journal of the Acoustical Society of America},
  137(2):822--831, 2015.

\bibitem{ishi2006acoustic}
Carlos Ishi, Hiroshi Ishiguro, and Norihiro Hagita.
\newblock Acoustic analysis of pressed voice.
\newblock {\em The Journal of the Acoustical Society of America},
  120(5):3374--3374, 2006.

\bibitem{ncvs.org}
Sources of vocal fluctuations.

\bibitem{hanson1999glottal}
Helen~M Hanson and Erika~S Chuang.
\newblock Glottal characteristics of male speakers: Acoustic correlates and
  comparison with female data.
\newblock {\em The Journal of the Acoustical Society of America},
  106(2):1064--1077, 1999.

\bibitem{gobl1988b}
Christer Gobl and Ailbhe~Nf Chasaide.
\newblock B. the effects of adjacent voiced/voice-less consonants on the vowel
  voice source: A cross language study.
\newblock 1988.

\bibitem{gobl1989preliminary}
Christer Gobl.
\newblock A preliminary study of acoustic voice quality correlates.
\newblock {\em STL-QPSR}, 4:9--21, 1989.

\bibitem{yumoto1982harmonics}
Eiji Yumoto, Wilbur~J Gould, and Thomas Baer.
\newblock Harmonics-to-noise ratio as an index of the degree of hoarseness.
\newblock {\em The journal of the Acoustical Society of America},
  71(6):1544--1550, 1982.

\bibitem{pinto1990unification}
Neal~B Pinto and Ingo~R Titze.
\newblock Unification of perturbation measures in speech signals.
\newblock {\em The Journal of the Acoustical Society of America},
  87(3):1278--1289, 1990.

\bibitem{kasuya1986acoustic}
Hideki Kasuya, Shigeki Ogawa, Yoshinobu Kikuchi, and Satoshi Ebihara.
\newblock An acoustic analysis of pathological voice and its application to the
  evaluation of laryngeal pathology.
\newblock {\em Speech Communication}, 5(2):171--181, 1986.

\bibitem{hollien1973method}
Harry Hollien.
\newblock A method for analyzing vocal jitter in sustained phonation.
\newblock {\em J. Phonetics}, 1:85--91, 1973.

\bibitem{koike1973application}
Yasuo Koike.
\newblock Application of some acoustic measures for the evaluation of laryngeal
  dysfunction.
\newblock 1973.

\bibitem{kataoka1996spectral}
Ryuta Kataoka, Ken-Ichi Michi, Kaoru Okabe, Tanetoshi Miura, and Hiroshi
  Yoshida.
\newblock Spectral properties and quantitative evaluation of hypernasality in
  vowels.
\newblock {\em The Cleft palate-craniofacial journal}, 33(1):43--50, 1996.

\bibitem{kataoka2001relationship}
Ryuta Kataoka, Donald~W Warren, David~J Zajac, Robert Mayo, and Richard~W Lutz.
\newblock The relationship between spectral characteristics and perceived
  hypernasality in children.
\newblock {\em The Journal of the Acoustical Society of America},
  109(5):2181--2189, 2001.

\bibitem{yoshida2000spectral}
H~Yoshida, Y~Furuya, K~Shimodaira, T~Kanazawa, R~Kataoka, and K~Takahashi.
\newblock Spectral characteristics of hypernasality in maxillectomy patients 1.
\newblock {\em Journal of Oral Rehabilitation}, 27(8):723--730, 2000.

\bibitem{lee2003acoustic}
Alice S-Y Lee, Valter Ciocca, and Tara~L Whitehill.
\newblock Acoustic correlates of hypernasality.
\newblock {\em Clinical Linguistics \& Phonetics}, 17(4-5):259--264, 2003.

\bibitem{hollien1972vocal}
Harry Hollien.
\newblock On vocal registers.
\newblock 1972.

\bibitem{titze1998principles}
Ingo~R Titze and Daniel~W Martin.
\newblock Principles of voice production, 1998.

\bibitem{titze2001acoustic}
Ingo~R Titze.
\newblock Acoustic interpretation of resonant voice.
\newblock {\em Journal of voice}, 15(4):519--528, 2001.

\bibitem{barrichelo2009resonant}
Viviane Barrichelo-Lindstr{\"o}m and Mara Behlau.
\newblock Resonant voice in acting students: perceptual and acoustic correlates
  of the trained y-buzz by lessac.
\newblock {\em Journal of Voice}, 23(5):603--609, 2009.

\bibitem{fraile2014cepstral}
Rub{\'e}n Fraile and Juan~Ignacio Godino-Llorente.
\newblock Cepstral peak prominence: A comprehensive analysis.
\newblock {\em Biomedical Signal Processing and Control}, 14:42--54, 2014.

\bibitem{lowell2012spectral}
Soren~Y Lowell, Richard~T Kelley, Shaheen~N Awan, Raymond~H Colton, and
  Natalie~H Chan.
\newblock Spectral-and cepstral-based acoustic features of dysphonic, strained
  voice quality.
\newblock {\em Annals of Otology, Rhinology \& Laryngology}, 121(8):539--548,
  2012.

\bibitem{lowell2011spectral}
Soren~Y Lowell, Raymond~H Colton, Richard~T Kelley, and Youngmee~C Hahn.
\newblock Spectral-and cepstral-based measures during continuous speech:
  capacity to distinguish dysphonia and consistency within a speaker.
\newblock {\em Journal of Voice}, 25(5):e223--e232, 2011.

\bibitem{shao2010acoustic}
Jun Shao, Julia~K MacCallum, Yu~Zhang, Alicia Sprecher, and Jack~J Jiang.
\newblock Acoustic analysis of the tremulous voice: assessing the utility of
  the correlation dimension and perturbation parameters.
\newblock {\em Journal of communication disorders}, 43(1):35--44, 2010.

\bibitem{story2001relationship}
Brad~H Story, Ingo~R Titze, and Eric~A Hoffman.
\newblock The relationship of vocal tract shape to three voice qualities.
\newblock {\em The Journal of the Acoustical Society of America},
  109(4):1651--1667, 2001.

\bibitem{bowler1964fundamental}
Ned~W Bowler.
\newblock A fundamental frequency analysis of harsh vocal quality.
\newblock {\em Communications Monographs}, 31(2):128--134, 1964.

\bibitem{scherer1973voice}
Klaus~R Scherer, Harvey London, and Jared~J Wolf.
\newblock The voice of confidence: Paralinguistic cues and audience evaluation.
\newblock {\em Journal of Research in Personality}, 7(1):31--44, 1973.

\bibitem{gallardo2017perceived}
Laura~Fern{\'a}ndez Gallardo and Benjamin Weiss.
\newblock Perceived interpersonal speaker attributes and their acoustic
  features.
\newblock {\em Preface \& Acknowledgements}, page~61, 2017.

\bibitem{eyben2016geneva}
Florian Eyben, Klaus~R Scherer, Bj{\"o}rn~W Schuller, Johan Sundberg, Elisabeth
  Andr{\'e}, Carlos Busso, Laurence~Y Devillers, Julien Epps, Petri Laukka,
  Shrikanth~S Narayanan, et~al.
\newblock The geneva minimalistic acoustic parameter set (gemaps) for voice
  research and affective computing.
\newblock {\em IEEE Transactions on Affective Computing}, 7(2):190--202, 2016.

\bibitem{eadie2005classification}
Tanya~L Eadie and Philip~C Doyle.
\newblock Classification of dysphonic voice: acoustic and auditory-perceptual
  measures.
\newblock {\em Journal of Voice}, 19(1):1--14, 2005.

\bibitem{heman2002relationship}
Yolanda~D Heman-Ackah, Deirdre~D Michael, and George~S Goding~Jr.
\newblock The relationship between cepstral peak prominence and selected
  parameters of dysphonia.
\newblock {\em Journal of Voice}, 16(1):20--27, 2002.

\bibitem{heman2003cepstral}
Yolanda~D Heman-Ackah, Deirdre~D Michael, Margaret~M Baroody, Rosemary
  Ostrowski, James Hillenbrand, Reinhardt~J Heuer, Michelle Horman, and
  Robert~T Sataloff.
\newblock Cepstral peak prominence: a more reliable measure of dysphonia.
\newblock {\em Annals of Otology, Rhinology \& Laryngology}, 112(4):324--333,
  2003.

\bibitem{hildebrand2016creaky}
Nicole Hildebrand-Edgar.
\newblock {\em Creaky voice: an interactional resource for indexing authority}.
\newblock PhD thesis, 2016.

\bibitem{steinmetz_2011}
Katy Steinmetz.
\newblock Katy steinmetz.
\newblock {\em Get Your Creak On: Is ‘Vocal Fry’ a Female Fad?}, Dec 2011.

\bibitem{yuasa2010creaky}
Ikuko~Patricia Yuasa.
\newblock Creaky voice: A new feminine voice quality for young urban-oriented
  upwardly mobile american women?
\newblock {\em American Speech}, 85(3):315--337, 2010.

\bibitem{weiss2010voice}
Benjamin Weiss and Felix Burkhardt.
\newblock Voice attributes affecting likability perception.
\newblock In {\em Eleventh Annual Conference of the International Speech
  Communication Association}, 2010.

\bibitem{hodges2010different}
Carolyn~R Hodges-Simeon, Steven~JC Gaulin, and David~A Puts.
\newblock Different vocal parameters predict perceptions of dominance and
  attractiveness.
\newblock {\em Human Nature}, 21(4):406--427, 2010.

\end{thebibliography}

\documentclass{article}

\usepackage[table]{xcolor}
\usepackage[margin=1in]{geometry}
\usepackage{tabularx}
\usepackage{enumitem}
\usepackage{longtable,tabu}
\setlist{nolistsep}
\usepackage{array,booktabs}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      

\usepackage{multirow}
\usepackage{booktabs}

\usepackage{longtable}
\usepackage{threeparttable}
\usepackage{textcomp}
\usepackage{array}

\usepackage{lipsum}

\newcommand\blfootnote[1]{%
  \begingroup
  \renewcommand\thefootnote{}\footnote{#1}%
  \addtocounter{footnote}{-1}%
  \endgroup
}

\title{Acoustic Correlates of the Voice Qualifiers: A Survey\blfootnote{Preprint. Work in progress.}}

\author{
  Shahan Ali Memon\\
  Language Technologies Institute \\
  Carnegie Mellon University \\
  Pittsburgh, PA 15213 \\
  \texttt{\{samemon\}}@cs.cmu.edu
}
\date{}
\begin{document}

\maketitle

\begin{abstract}
Our voices are as distinctive as our faces and fingerprints. There is a spectrum of non-disjoint traits that make our voices unique and identifiable, such as the fundamental frequency, the intensity, and most interestingly the quality of the speech. Voice quality refers to the characteristic features of an individual's voice.

Previous research has from time-to-time proven the ubiquity of voice quality in making different paralinguistic inferences. These inferences range from identifying personality traits, to health conditions and beyond.

In that context, one could arguably summarize the constituent perceptual elements of voice quality into two categories: primary qualities, and voice qualifiers. Primary qualities, as defined by Poyatos \cite{poyatos1991paralinguistic}, are well-defined traits that are always present in the human voice. These include timbre, loudness, pitch, intonation range, etc.

However, what makes our voices truly unique is a spectrum of complex \emph{voice qualifiers} which differ significantly across different speakers and may in fact define different speaker states and traits.

In loose terms, Dr. Ingo Titze defines these voice qualifiers as proxy features in determination of ones' genetic and anatomical structure, and the learned components. In the 8th Vocal Fold Physiology Conference in April 1994, he proposed a sufficiently comprehensive list of voice qualifiers which we use as basis for our research.

In this manuscript, we first map the paralinguistic voice qualifiers to their acoustic correlates in the light of the previous research and literature. We also determine the openSMILE correlates one could possibly use to measure those correlates. In the second part, we give a set of example paralinguistic inferences that can be made using different acoustic and perceptual voice quality features.

\end{abstract}
\pagebreak
\section{Acoustic Correlates of Voice Quality}
\label{sec:vqtable}
Table~\ref{tab:1} maps each of the voice quality features to its definition, and the acoustic correlates found in the background literature found as useful cues to measure it. We also provide the possible openSMILE\cite{eyben2010opensmile} features that could be used as proxy features in relation to the acoustic features. 
\begin{longtable}{
p{1.5cm}
p{2.0cm}
p{2.0cm}
p{4.0cm}
p{5.0cm}
}
\caption{Acoustic Correlates of Voice Quality}\label{tab:1}\\
 \toprule  
\multicolumn{1}{c}{Voice Quality}&
\multicolumn{1}{c}{Perceptive Definition}&
\multicolumn{1}{c}{Physiology}&
\multicolumn{1}{c}{Correlate(s)}&
\multicolumn{1}{c}{openSMILE Correlate(s)}\\
\midrule
\endfirsthead
\caption[]{Acoustic Correlates of Voice Quality}\\
 \toprule  
\multicolumn{1}{c}{Voice Quality}&
\multicolumn{1}{c}{Perceptive Definition}&
\multicolumn{1}{c}{Physiology}&
\multicolumn{1}{c}{Correlate(s)}&
\multicolumn{1}{c}{openSMILE Correlate(s)}\\ \midrule
\endhead

Aphonic & no sound or whisper & inability to set vocal folds into vibration, caused by lack of appropriate power (air pressure) or a muscular/tissue problem of the folds & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Pitch Absent
    \item Lower HNR
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_amean
    \item logHNR\_sma\_amean
\end{itemize}
\\
\hline
Biphonic & two independent pitches & two sources of sound (e.g., true folds and false folds, or two folds and whistle due to vortex in air) &

\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F1 within 10\% of pitch
    \item Roughness \cite{keating2015acoustic}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
	\item F0final\_sma\_amean
    \item F1frequency\_sma3nz\_amean
    \item openSMILE correlates for Roughness from ~\ref{tab:1}
\end{itemize}
\\
\hline
Breathy & sound of air is apparent or impression of glottal air leakage and turbulence noise during phonation & noise is caused by turbulence in or near glottis, caused by loose valving of laryngeal muscles (lateral cricoarytenoid, interarytenoid and posterior cricoarytenoid) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Pitch Amplitude (PA) - supposed to be lower for voiced sounds since PA represents the degree of voicing \cite{doi:10.1044/jshr.3302.298}
    \item \%Jitter which is the mean jitter in ms divided by the mean period in ms - found to be higher for breathy voices. \cite{doi:10.1044/jshr.3302.298}
    \item High Peak levels of the first harmonic \cite{de1995some}
    \item Slope of log-magnitude spectrum in 2-5 KHz band; the steepness of the spectral slope is strongly related to the glottal closure; the faster the glottis is closed, the more pulse like excitation resulting in flat harmonic source spectrum. Breathy voices are gradual - with insufficient glottal closure, yielding spectral slope more than 12 dB per octave. \cite{de1995some}\cite{childers1991vocal}
    \item HNR in b1 and b2 \cite{de1995some}\cite{wolfe1987prediction}
    \item F0 estimated in the cepstral domain \cite{stranik2014acoustic}
    \item glottal-to-noise excitation ratio
\cite{stranik2014acoustic}
	\item high-to-mid or low-frequency energy
\cite{stranik2014acoustic}
	\item ratio of number of voiced parts to number of all samples in the utterance \cite{stranik2014acoustic}
    \item Energy in 5-8 KHz band $\equiv$ Energy in 2-5 KHz band for breathy utterances; non-breathy utterances exhibit lower energy in 5-8 KHz region. \cite{labuschagne2016perception}\cite{hammarberg1980perceptual}
    \item H1-H2 \cite{hillenbrand1994acoustic}\cite{labuschagne2016perception}\cite{klatt1990analysis}
    \item Higher $\frac{H}{L}$ found for breathy voices \cite{hillenbrand1996acoustic}
  \end{itemize} &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item jitterLocal\_sma\_amean
    \item logHNR\_sma\_amean
    \item F0semitoneFrom27.5Hz
    \item F0final\_sma\_amean
    \item logRelF0-H1-H2\_sma3nz\_amean
    \item voicingFinalUnclipped
    \item pcm\_RMSenergy
\end{itemize}\\
  \\

\hline
Covered & muffled or 'darkened' sound  & lips are rounded and protruded or larynx is lowered to lower all formants so a stronger fundamental is obtained &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item High Spectral Entropy
    \item Lower F1,F2, and F3
    \item Stronger F0
\end{itemize} &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item pcm\_fftMag\_spectralEntropy
    \item F1frequency\_sma3nz\_amean
    \item F2frequency\_sma3nz\_amean
    \item F3frequency\_sma3nz\_amean
    \item F0final\_sma3nz\_amean
\end{itemize}\\
\hline
Creaky & sounding like two hard surfaces rubbing against one another & a complex pattern of vibrations in the vocal folds creates a intricate formation of subharmonics and modulations & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower F0 \cite{garellek2012timing}\cite{gerfen2005production}
    \item irregular F0 \cite{keating2015acoustic}
    \item higher values of SHR (subharmonics-to-harmonic ratio) \cite{garellek2015perception}
    \item relatively weak H1 (due to low airflow through constricted glottis)
    \item low B1 values \cite{garellek2015perception}
    \item relatively strong frequency harmonics (which together result in lower values of various harmonic difference measures (H1-A1, H1-A2, H1-A3, H2-H4, H4-2k, 2K-5k) \cite{keating2015acoustic}
    \item  low H1-H2 \cite{garellek2015perception}
    \item less spectral tilt (low flow through glottis means less energy in H1, So harmonic amplitude difference have lower values hence less spectral tilt) \cite{garellek2015perception}
\end{itemize} &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_amean
    \item jitterLocal\_sma\_amean
    \item logRelF0-H1-H2\_sma3nz\_amean
    \item logRelF0-H1-A3\_sma3nz\_amean
    \item pcm\_fftMag\_mfcc\_sma[0]
    \item pcm\_fftMag\_mfcc\_sma[1]
    \item pcm\_fftMag\_mfcc\_sma[2]
    \item pcm\_fftMag\_mfcc\_sma[3]
\end{itemize}\\
\hline
Diplophonic & pitch supplemented with another pitch one octave lower, roughness usually apparent & a period doubling, or F0/2 subharmonic & Non-pressed creaky voice \cite{ishi2006acoustic} & openSMILE correlates for pressed and creaky voice in ~\ref{tab:1}\\
\hline
Flutter & often called bleat because it sounds like a lamb's cry & amplitude changes or frequency modulations in the 8-12Hz range &  
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Percent Jitter \cite{ncvs.org}
    \item Percent Shimmer \cite{ncvs.org}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
	\item jitterLocal\_sma\_amean
    \item shimmerLocal\_sma\_amean
\end{itemize}
\\
\hline
Glottalized & clicking noise heard during voicing & forceful adduction or abduction of the vocal folds during speech & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Spectral Tilt \cite{hanson1999glottal}
    \item Increased Downward Spectral Slope \cite{gobl1988b}\cite{gobl1989preliminary}
    \item Decreased F1 Amplitude \cite{gobl1988b}\cite{gobl1989preliminary}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
	\item pcm\_fftMag\_mfcc\_sma[0]
    \item pcm\_fftMag\_mfcc\_sma[1]
    \item pcm\_fftMag\_mfcc\_sma[2]
    \item pcm\_fftMag\_mfcc\_sma[3]
	\item F1amplitudeLogRelF0
    \item pcm\_fftMag\_spectralSlope
\end{itemize}
\\
\\
\hline
Raspy or Hoarse & harsh, grating sound & combination of irregularity in vocal fold vibration and glottal noise generation &  
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Pitch Amplitude (PA) - supposed to be lower for hoarse voices \cite{doi:10.1044/jshr.3302.298}
    \item Lower HNR \cite{doi:10.1044/jshr.3302.298} \cite{yumoto1982harmonics}
    \item Jitter and Shimmer \cite{pinto1990unification}\cite{kasuya1986acoustic}\cite{hollien1973method}\cite{koike1973application}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
	\item logHNR\_sma\_amean
	\item jitterLocal\_sma\_amean
    \item shimmerLocal\_sma\_amean
\end{itemize}
\\
\hline
Nasal or Honky & excessive nasality & excessive acoustic energy couples to the nasal tract & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Increase in power level between the first and second formant. \cite{kataoka1996spectral}\cite{kataoka2001relationship}\cite{yoshida2000spectral}
    \item Decrease in power level in second and third formants region. \cite{kataoka1996spectral}\cite{kataoka2001relationship}\cite{yoshida2000spectral}
    \item Higher amplitudes for the bands centered at 630, 800 and 1000 Hz. \cite{lee2003acoustic}
    \item Significantly lower amplitude for the band centered at 2500 Hz. \cite{lee2003acoustic}
    \item One-third-octave analysis. \cite{kataoka2001relationship}\cite{kataoka1996spectral}
    \item Decrease in the amplitude of the first formant (F1) - Reduced F1 amplitude is associated with the increase in formant bandwidths and upward shifts in formant frequencies. \cite{lee2003acoustic}
    
\end{itemize} &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F1amplitudeLogRelF0
    \item F2amplitudeLogRelF0
   	\item F3amplitudeLogRelF0
\end{itemize}\\
\\
\hline
Jitter & pitch sounds rough & fundamental frequency varies from cycle to cycle &  Jitter/Jita & jitterLocal\_sma\_amean \\ 
\hline
Pressed & harsh, often loud (strident) quality & vocal processes of the arytenoid cartilages are squeezed together, constricting the glottis, and causing low airflow and medial compression of the vocal folds &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0 not low \cite{keating2015acoustic}
    \item Glottal constriction (low H1-H2) \cite{keating2015acoustic}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item 
\end{itemize}
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_amean
    \item logRelF0-H1-H2\_sma3nz\_amean
\end{itemize}
\\
\hline
Pulsed or Strohbass or Vocal Fry & sounds similar to food cooking in a hot frying pan & sound gaps caused by intermittent energy packets below 70 Hz and formant energy dies out prior to re-excitation & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower Pitch Amplitude (PA) \cite{doi:10.1044/jshr.3302.298}
    \item Lower HNR \cite{doi:10.1044/jshr.3302.298}\cite{keating2015acoustic}
    \item Low F0 \cite{keating2015acoustic}\cite{hollien1972vocal}
    \item F0 less than 70 Hz. \cite{titze1998principles}
    \ Glottal constriction (low H1-H2) \cite{keating2015acoustic}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item logHNR\_sma\_amean
    \item F0final\_sma\_amean
    \item logRelF0-H1-H2\_sma3nz\_amean
\end{itemize}
\\
\hline
Resonant & brightened or 'ringing' sound that carries well & epilaryngeal resonance is enhanced, producing a strong spectral peak at 2500-3500 Hz; in effect, formants F3, F4 and F5 are clustered &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Not breathy \cite{titze2001acoustic} 
    \item Not pressed \cite{titze2001acoustic} 
    \item Low formant frequencies \cite{barrichelo2009resonant}
    \item Significant difference between F1 and H2 for the male \cite{barrichelo2009resonant}
    \item Significantly small difference between F1 and F0 for the female \cite{barrichelo2009resonant}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item openSMILE correlates of breathy voice from ~\ref{tab:1} 
    \item openSMILE correlates of pressed voice from ~\ref{tab:1} 
    \item F1frequency\_sma3nz\_amean
    \item F2frequency\_sma3nz\_amean
    \item F3frequency\_sma3nz\_amean
\end{itemize}
\\
\hline
Rough & uneven, bumpy sound appearing to be unsteady short-term, but persisting over the long-term or presence of a low-frequency noise component & modes of vibration of the vocal folds are not synchronized & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Pitch Amplitude (PA) \cite{doi:10.1044/jshr.3302.298}
    \item Lower HNR \cite{doi:10.1044/jshr.3302.298}
    \item Lower SFR (the flatness of magnitude spectrum (in dB) is the log of the ratio of the Geometric mean of the spectrum to Arithmetic mean of the spectrum \cite{doi:10.1044/jshr.3302.298}
    \item Mean HNR in b1 (400-2KHz) and b2 (2KHz-5KHz) bands \cite{de1995some}\cite{wolfe1987prediction}\cite{yumoto1982harmonics}
\end{itemize} &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item logHNR\_sma\_amean
    \item audSpec\_Rfilt\_sma\_flatness
\end{itemize}
\\
\hline
Shimmer & crackly, buzzy & short-term (cycle-to-cycle) variation in a signal's amplitude & shimmer & shimmerLocal\_sma\_amean \\
\hline
Strained & effortfulness apparent in voice, hyperfunction of neck muscles, entire larynx may compress & excessive energy focused in laryngeal region & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Cepstral peak prominence (CPP) \cite{fraile2014cepstral}\cite{lowell2012spectral}
    \item L/H ratio (spectral) and its standard deviation\cite{lowell2012spectral}
    \item Spectral Skewness/Tilt \cite{lowell2011spectral}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item pcm\_fftMag\_spectralSkewness
    \item any energy based feature for computing L/H (e.g. pcm\_RMSenergy)
\end{itemize} \\
\hline
Tremerous & affected by trembling or tremors & modulation of 1-15 Hz in either amplitude or pitch due to a neurological or biomechanical cause &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Percent Jitter \cite{shao2010acoustic}
    \item Percent Shimmer \cite{shao2010acoustic}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
	\item jitterLocal\_sma\_amean
    \item shimmerLocal\_sma\_amean
\end{itemize}
\\
\hline
Twangy & sharp, bright sound & often attributed to excessive nasality, but probably also has an epilaryngeal basis &  F1 and F2 tend to be spread far apart for twangy vowels \cite{story2001relationship} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F1Frequency\_sma3nz\_amean 
    \item F2Frequency\_sma3nz\_amean
    \item F1bandwidth\_sma3nz\_amean
    \item F2bandwidth\_sma3nz\_amean
\end{itemize}\\
\hline
Ventricular or harsh & very rough (Louis Armstrong-type voice) & phonation using the false folds anterior rather than the vocal folds; unless intentional due to damage to the true folds, considered an abnormal muscle pattern dysphonia & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0 range twice that of the range for the non-harsh voice \cite{bowler1964fundamental}
    \item Distribution of F0 markedly skewed toward the lower end of the distribution \cite{bowler1964fundamental}
    \item Abrupt and extreme changes in the fundamental frequency (Jitter) \cite{bowler1964fundamental}
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item F0final\_sma\_peakDistStddev
    \item jitterLocal\_sma\_amean
\end{itemize}\\
\\
\hline
Wobble & wavering or irregular variation in sound & amplitude and/or frequency modulations in the 1-3 Hz range &  FM and/or AM $\leq$ 3 Hz &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item jitterLocal\_sma\_amean
    \item shimmerLocal\_sma\_amean
\end{itemize}\\
\\
\hline
Yawny & quality is akin to sounds made during a yawn & larynx is lowered and pharynx is widened, as people do when yawning - hence the name & formants F1 and F2 in the yawn vowel quality tend to be more closely spaced than in the normal vowels \cite{story2001relationship} &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F1Frequency\_sma3nz\_amean 
    \item F2Frequency\_sma3nz\_amean
    \item F1bandwidth\_sma3nz\_amean
    \item F2bandwidth\_sma3nz\_amean
\end{itemize}\\
\bottomrule
\end{longtable}

\section{Inference of Speaker Traits and States from Acoustic Cues}
\label{sec:paralinguistictable}
Table~\ref{tab:2} defines the mapping from acoustic correlates or voice qualifiers to paralinguistic inferences researchers have made over the years across different literature.
\begin{longtable}{
p{1cm}
p{5.5cm}
p{7cm}
p{1cm}
}
\caption{Paralinguistic Inferences from Voice}\label{tab:2}\\
 \toprule  
\multicolumn{1}{c}{Meta-Inference}&
\multicolumn{1}{c}{Acoustic or Perceptual Correlate(s)}&
\multicolumn{1}{c}{openSMILE Correlate(s)}&
\multicolumn{1}{c}{Source}\\
\midrule
\endfirsthead
\caption[]{Paralinguistic Inferences from Voice}\\
 \toprule  
\multicolumn{1}{c}{Meta-Inference}&
\multicolumn{1}{c}{Acoustic or Perceptual Correlate(s)}&
\multicolumn{1}{c}{openSMILE Correlate(s)}&
\multicolumn{1}{c}{Source}\\
\midrule
\endhead

Confidence & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Greater spectral energy 
    \item Greater Pitch
    \item Faster rate of speech
    \item Less number of pauses
    \item Less duration of pauses in comparison to doubtful voice
\end{itemize} & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item pcm\_RMSenergy 
    \item pcm\_zcr
    \item F0
\end{itemize}
& \cite{scherer1973voice} \\
\hline
Enthusiastic, Forceful, Active, Competent & Confidence &
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item pcm\_RMSenergy 
    \item pcm\_zcr
    \item F0
\end{itemize}
& \cite{scherer1973voice} \\
\hline
Doubtful & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Less spectral energy 
    \item Lower Pitch
    \item Slower rate of speech
    \item More number of pauses
    \item Longer pauses 
\end{itemize}
& 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item pcm\_RMSenergy 
    \item pcm\_zcr
    \item F0
\end{itemize}
& \cite{scherer1973voice} \\
\hline
Warmth (Male) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Higher F0 range 
    \item Higher Mean Spectral Slope (0-500Hz)
    \item Lower Standard Deviation (SD) of F1 and F2
\end{itemize}
& 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item pcm\_fftMag\_spectralSlope
    \item slopeV0\-500\_sma3nz\_amean
    \item slopeUV0\-500\_sma3nz\_amean
    \item F1frequency\_sma3nz\_stddevNorm
    \item F2frequency\_sma3nz\_stddevNorm
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Attractiveness (Male) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Higher F0 range 
    \item Higher Standard Deviation (SD) of Hammarberg Index
    \item Lower mean length of unvoiced segments
    \item Higher SD of F0
    \item Lower median of F0
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item hammarbergIndexV\_sma3nz\_stddevNorm
    \item hammarbergIndexUV\_sma3nz\_stddevNorm
    \item voicingFinalUnclipped
    \item F0final\_sma\_stddev
    \item F0semitoneFrom27.5Hz\_sma3nz\_stddevNorm
    \item F0semitoneFrom27.5Hz\_sma3nz\_percentile50.0
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Confidence (Male) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower median F0 
    \item Higher F0 range
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item F0semitoneFrom27.5Hz\_sma3nz\_percentile50.0
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Compliance (Male) & 
Lower SD length of voiced segments
&
voicingFinalUnclipped
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Maturity (Male) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower median F0 
    \item Higher SD of F3
    \item Higher SD of F3 bandwidth
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0semitoneFrom27.5Hz
    \item F3frequency\_sma3nz\_stddevNorm
    \item F3bandwidth\_sma3nz\_stddevNorm
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Warmth (Female) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Higher F0 range
    \item Higher F1
    \item Higher SD of F2 bandwidth
    \item Higher SD of spectral flux
    \item Lower SD of F1
    \item Higher median F0
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item F1frequency\_sma3nz\_amean
    \item F2bandwidth\_sma3nz\_stddevNorm
    \item pcm\_fftMag\_spectralFlux
    \item F1frequency\_sma3nz\_stddevNorm
    \item F0semitoneFrom27.5Hz
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Attractiveness (Female) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]	
	\item Higher F0 range
    \item Higher F1
    \item Higher SD of F2 bandwidth
    \item Lower SD of F1
    \item Lower mean spectral slope (0-500 Hz)
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item F1frequency\_sma3nz\_amean
    \item F2bandwidth\_sma3nz\_stddevNorm
    \item F1frequency\_sma3nz\_stddevNorm
    \item slopeV0\-500\_sma3nz\_amean
    \item slopeUV0\-500\_sma3nz\_amean
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Confidence (Female) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
	\item Higher F0 range
    \item Higher SD of falling slope of loudness
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0final\_sma\_range
    \item loudness\_sma3\_stddevFallingSlope
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Compliance (Female) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower SD of F1
    \item Higher F1
    \item Lower loudness range
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F1frequency\_sma3nz\_stddevNorm
    \item F1frequency\_sma3nz\_amean
    \item loudness\_sma3\_amean
    \item loudness\_sma3\_stddevNorm
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Maturity (Female) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower median F0
    \item Higher mean mfcc4
    \item Lower F1
    \item Higher mean mfcc2
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0semitoneFrom27.5Hz or F0final
    \item mfcc4\_sma3\_amean
    \item F1frequency\_sma3nz\_amean
    \item mfcc2\_sma3\_amean
\end{itemize}
& \cite{gallardo2017perceived}\cite{eyben2016geneva} \\
\hline
Dysphonia & 
Strained Voice Quality
&
Acoustic correlates of strained voice quality from ~\ref{tab:1}
& \cite{eadie2005classification}\cite{heman2002relationship}\cite{heman2003cepstral} \\
\hline
Vocal Tremor & 
Tremerous Voice Quality
&
Acoustic correlates of tremerous voice quality from ~\ref{tab:1}
& \cite{garellek2012timing} \\
\hline
Authoritative Personality & 
Confidence, Creakiness
&
Acoustic correlates of confidence, and creaky voice quality from ~\ref{tab:2} and ~\ref{tab:1} respectively
& \cite{hildebrand2016creaky}\cite{steinmetz_2011}\cite{yuasa2010creaky} \\
\hline
Likability (male) & 
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item Lower F0
    \item s\_cog (center of gravity for spectrum averaged for each stimulus)
\end{itemize}
&
\begin{itemize}[label={--},noitemsep,leftmargin=*,topsep=0pt,partopsep=0pt]
    \item F0semitoneFrom27.5Hz or F0final
    \item audSpec\_Rfilt\_sma\_centroid
\end{itemize}
& \cite{weiss2010voice} \\
\hline
Likability (female) & 
higher s\_sd (standard deviation for spectrum averaged for each stimulus)
&
spectralFlux\_sma3\_amean
& \cite{weiss2010voice} \\
\hline
Dominance (male) & 
Lower F0 variation
&
F0semitoneFrom27.5Hz or F0final
& \cite{hodges2010different} \\
\hline
Happiness, Joy, Confidence, Anger, Fear & 
High Pitch levels
&
F0\_mean
& \cite{hodges2010different} \\
\hline
\bottomrule
\end{longtable}

\bibliography{ref}
\bibliographystyle{unsrt}

\end{document}